### v2

# “We Demand Justice!”: Towards Social Context Grounding of Political Texts

Abstract
Social media discourse frequently consists of ‘seemingly similar language used by opposing sides of the political spectrum’, often translating to starkly contrasting perspectives. E.g., “thoughts and prayers”, could express sympathy for mass-shooting victims, or criticize the lack of legislative action on the issue. This paper defines the context required to fully understand such ambiguous statements in a computational setting and ground them in real-world entities, actions, and attitudes. We propose two challenging datasets that require an understanding of the real-world context of the text. We benchmark these datasets against models built upon large pre-trained models, such as RoBERTa and GPT-3. Additionally, we develop and benchmark more structured models building upon existing Discourse Contextualization Framework and Political Actor Representation models. We analyze the datasets and the predictions to obtain further insights into the pragmatic language understanding challenges posed by the proposed social grounding tasks.
“We Demand Justice!”: Towards Social Context Grounding of Political Texts
Rajkumar Pujari Purdue University, USA rpujari@purdue.edu                        Chengfei Wu Purdue University, USA wu1491@purdue.edu                        Dan Goldwasser Purdue University, USA dgoldwas@purdue.edu
1 Introduction
Over the past decade micro-blogging websites have become the primary channel of engagement between politicians and people in the US, resulting in an increasingly succinct political discourse. Politicians from the same party often coordinate the phrasing of social messaging in developing events to amplify their impact (Vaes et al., 2011; Weber and Neumann, 2021). Hence, repetitive phrases such as ‘Thoughts and Prayers’ are extensively used to signal more nuanced stances across several events. The interaction among politicians from opposing parties often leads to messaging phrased similarly but intended to signal opposing real-world actions. ‘Thoughts and Prayers’ is used to express condolences by Republicans in mass shooting events, and by Democrats convey an angry or sarcastic tone as a call for action demanding ‘tighter gun control measures’. The example in fig. 1 shows contrasting interpretations of the phrase “We need to keep our teachers safe!” depending on different speakers and in the context of different events.
Figure 1: An example of varied intended meanings behind the same political message depending on the Author and Event in context
Tweet Target Entity and Sentiment Vague Text Disambiguation
Tweet: As if we needed more evidence. #kavanaugh
Vague Text: First, but not the last.
Event: Kavanaugh Supreme Court Nomination
Event: US withdraws from Paris climate agreement that
enforces environmental targets after three years
Author: Earl Blumenauer (Democrat Politician) Author Party: Republican
Targets: Brett Kavanaugh (negative), Julie Swetnick (positive)
Christine Ford (positive), Deborah Ramirez (positive)
Disambiguation: The withdrawal from the Paris climate agreement
is the first step of many to come for the Trump administration. It will
not be the last, as more positive changes are sure to follow.
Incorrect Disambiguations: 1) Joe Biden’s inauguration marks the first day of a new era of progress and prosperity, lasting positive changes are coming. (Incorrect Event) 2) The Paris Climate Agreement withdrawal is the first of many backward steps this Trump administration is sure to take in destroying our environment. (Incorrect Stance) 3) This is the time for America to move forward and make progress without being held back by a global agreement that doesn’t serve our interests. (Doesn’t match the vague text)
Target Task Data Statistics Vague Text Data Statistics
Unique Tweets $865865$ Unique Vague Texts $9393$
Positive Targets $15131513$ Positive Examples $739739$
Negative Targets $10851085$ Negative Examples $22172217$
Neutral Targets $784784$ Total Examples $29562956$
Non-Targets $25092509$ Number of Events $99$
Total Data Examples $58915891$ Hard Test Examples $180180$
Number of Events $33$
Table 1: Examples of Annotated Datasets and their statistics
Humans familiar with the stances of a politician and, possessing knowledge about the event from the news, can easily understand the intended meaning. But, computationally understanding the correct meaning of such language is challenging. Our main question in this paper is - Can an NLP model find the right meaning? From a linguistic perspective, we follow the distinction (Bach, 2008) between semantic interpretation (where the meaning is encoded in the utterance directly and does not change based on its external context), and pragmatic interpretation (that depends on extra-linguistic information). The latter has gathered significant interest in the NLP community recently (Bender and Koller, 2020; Bisk et al., 2020), focusing on language understanding, when grounded in an external context (Fried et al., 2023). To a large extent, the focus of such studies was on grounding language in a perceptual environment (e.g., image captioning (Andreas and Klein, 2016; Sharma et al., 2018; Alikhani et al., 2020), instruction following (Wang et al., 2016; Suhr et al., 2019; Lachmy et al., 2022), and game playing (Potts, 2012; Udagawa and Aizawa, 2019) tasks). Unlike these works, in this paper, we focus on grounding language in a social context, i.e., modeling the common ground (Clark and Brennan, 1991; Traum, 1994; Stalnaker, 2002) between the author and their social media followers that enables understanding what would otherwise be highly ambiguous utterances.
The Social Context Understanding, needed for building successful models for such tasks, can come from a wide variety of sources. The politician’s party affiliation, historical stances on the issue, social relationships, knowledge about the involved entities, and related prior and upcoming events, are a few examples of the relevant social context. In fig. 1 event #1#1\#1# 1, combining the event information (school shooting) with the speakers’ stances on gun control, would facilitate understanding the intended meaning of the text better.
This paper operationalizes the problem of ‘Social Context Grounding’ by defining tasks capturing the intended meaning at different level of granularity. At the most basic level, providing the social context should help identify the entities targeted and the sentiment towards them. In fig. 1, the social context ⟨Ev.#1, Harris⟩and the text “we need to keep our teachers safe” $\Rightarrow$ negative attitude towards guns. A more nuanced account of meaning, which we formulate as a separate task, captures the specific means in which the negative attitude is expressed (the Interpretation in fig. 1). We present two datasets corresponding to these tasks, namely, ‘Target Entity and Sentiment Detection’ and ‘Vague Text Disambiguation’. In the first, the goal is to predict: $11$) whether a given entity is the intended target of a politician’s tweet and $22$) the sentiment towards the intended targets. We explicitly focus on tweets that do not always mention the targets in their text to incentivize modeling the pragmatic communicative intent of the text. In the second task, given an ambiguous political message such as “We demand justice” and its social context (associated event, & the author’s party affiliation), the task is to identify a plausible unambiguous explanation of the message. Note that the ground truth for all these tasks is based on human pragmatic interpretation, i.e., “guns” is a negative target of “we need to keep our teachers safe”, despite not being mentioned in the text, since it was perceived in this way by a team of human annotators reading the tweet and knowing social context. We show examples of each task in table 1. We describe the datasets in detail in section 3.
We evaluate the performance of various models, as a way to test the need for social context and compare different approaches for modeling it. These include pre-trained LM-based classifiers, and LLM in-context learning (Brown et al., 2020a; Black et al., 2022), which use a textual representation of the social context. We adopt an existing graph-based discourse contextualization framework (Pujari and Goldwasser, 2021; Feng et al., 2022) to explicitly model the social context needed to solve the proposed tasks. Our results demonstrate that discourse contextualization models outperform other models on both tasks. We present an error analysis of the outputs to gain further insights. We describe the models in section 4 and the results in section 5.
We also present qualitative visualization of a political event, Brett Kavanaugh Supreme Court Nomination in section 6.4 from the lens of target entity-sentiment to showcase a unique summary of the event discourse. Then, we perform human evaluation on ‘Vague Text Disambiguation’ and observe that humans find the task much easier than the evaluated models. We also present observations of human vs. LLM errors in disambiguations. Our contributions can be summarized as:
  1. 1.
    Defining and operationalizing the ‘Social Context Grounding’ task in political discourse
  2. 2.
    Evaluating various state-of-the-art context representation models on the task. Adopting existing discourse contextualization framework for the proposed tasks and evaluating GPT-3 in-context learning performance on the tasks
  3. 3.
    Performing human studies to benchmark the dataset difficulty and GPT-3 generation performance comparison with human workers [^1] Our data and code will be released under MIT license
2 Related Work
Pragmatic Language Grounding gained significant focus recently (Bender and Koller, 2020; Bisk et al., 2020) following the rise of Pretrained Language Models (Devlin et al., 2019; Liu et al., 2019; Brown et al., 2020a) as unified NLP models. Most grounding tasks address multi-modal or physical environment descriptions (Barnard et al., 2003; Vogel and Jurafsky, 2010; Chen and Mooney, 2011; Tellex et al., 2011; Mitchell et al., 2012; Anderson et al., 2018). We refer the reader to Fried et al. (2023) for a thorough overview. In contrast, we focus on grounding language in a social context.
Social Context Modeling Hovy and Yang (2021) show that modeling social context is necessary for human-level NLU. As political messages are often targeted at the voter base aware of the political context (Weber and Neumann, 2021; Vaes et al., 2011), they are vague by design. Several previous works model social context for entity linking (Yang et al., 2016), social media connections relationship for fake news detection Baly et al. (2018); Mehta et al. (2022) and, political bias detection (Li and Goldwasser, 2019; Baly et al., 2020). These works model partial aspects of social context, relevant to their tasks. Two recent frameworks aim to capture social context holistically (Pujari and Goldwasser, 2021; Feng et al., 2022). Evaluation tasks presented in both works show interesting social context understanding but are not fully representative of the challenges of Social Context Grounding. Zhan et al. (2023) propose a dataset for dialogue understanding addressing general social commonsense.
Related Semantic and Pragmatic tasks closest to our Target Entity Sentiment Identification task is Stance Detection in social media Mohammad et al. (2016); AlDayel and Magdy (2020). To clarify our contribution, Mohammad et al. (2016), a popular SemEval task, looks at sentiment towards $55$ targets, while our data has $362362$ unique targets. Allaway and McKeown (2020) and Zhang et al. (2022) also propose stance datasets on tweets. But, they focus mainly on semantic understanding of text that allows them to predict agreement or disagreement with well-defined statements. Our Vague Text Disambiguation task is related to recent works that study implicit inferences Hoyle et al. (2023), and pragmatic understanding Hu et al. (2023). However, our tasks evaluate pragmatic understanding using an explicit context, absent in those tasks.
3 Social Context Grounding Tasks
We design and collect two datasets for Social Context Grounding evaluation, and define three pragmatic interpretation tasks. In the Tweet Target Entity and Sentiment dataset, we collect annotations of opinionated tweets from known politicians for their intended targets and sentiments towards them. We focus on three political events for this task. The dataset and its collection are described below in section 3.1. In the Vague Text Disambiguation Task, we collect plausible explanations of vague texts, given the social context, consisting of author affiliation and specific event. We focus on eight political events. This dataset is detailed in section 3.2. Examples and data statistics are shown in table 1.
3.1 Tweet Target Entity and Sentiment Task
In this task, given a tweet T, its context, and an entity E, the objective is to predict whether or not E is a target of T and the sentiment towards E. Political discourse often contains opinionated discourse about world events and social issues. We collect tweets that don’t directly mention the target entities. Thus, connecting the text with the event details and the author’s general perspectives is necessary to solve this task effectively. We pick the focal entities for the given event and let human annotators expand on that initial set, based on their interpretation of the contextualized text. A target entity is conceptualized as an entity present in the full intended interpretation of the tweet.
We focus our tweet collection on three recent divisive events: George Floyd Protests, 2021 US Capitol Attacks, and Brett Kavanaugh’s Supreme Court Nomination. We identify relevant participating entities for each of the three events. Examples of the involved entities for the event George Floyd Protests were George Floyd, United States Police, Derek Chauvin, Donald Trump, Joe Biden, United States Congress, Black people, Democratic Party, Republican Party, BLM, Antifa.
3.1.1 Target-Sentiment Data Collection
We filter 3,45434543,4543 , 454 tweets for the three events using hashtags, keyword-based querying, and the dates of the event-based filtering from the Congress Tweets repository corpus [^2] https://github.com/alexlitel/congresstweets. We collect a subset of 1,77917791,7791 , 779 tweets that contain media (images/video) to increase the chances of the tweet text not containing the target entity mentions. Then, we use $66$ in-house human annotators and Amazon Mechanical Turk (AMT) workers who are familiar with the event context for annotation. We ask them to annotate the targeted entities and sentiments towards the targets. The authors of this paper also participated in the annotation process. We provide them with entity options based on the event in the focus of the tweet. Annotators are allowed to add additional options if needed. We also ask the annotators to mark non-targets for each tweet. We instruct them to keep the non-targets as relevant to the event as possible to create harder negative examples. Each tweet is annotated by three annotators. We filter $865865$ unique tweets with 5,89158915,8915 , 891 annotations, with majority agreement on each tweet. All the AMT annotations were additionally verified by in-house annotators for correctness. AMT workers were paid USD $11$ per tweet. It took $33$ minutes on average for each assignment, resulting in an hourly pay of USD $2020$. We include screenshots of the collection task GUIs in the appendix. We split the train, and test sets by events, authors, and targets to incentivize testing the general social grounding capabilities of the models. The test set also consists of authors, targets, and events not seen in the training set. We use Capitol Riots event for the test set of Target Entity and Sentiment Task. We split the examples into 4,37043704,3704 , 370 train, $511511$ development, and 1,00910091,0091 , 009 test examples.
3.2 Vague Text Disambiguation Task
The task of Vague Text Disambiguation is designed to capture pragmatic interpretation at a finer-grained level. It can be viewed as a variant of the well known paraphrase task, adapted for the social context settings. The model is evaluated on its ability to identify plausible interpretations (i.e., a sentence explicitly describing the author’s intent) of an ambiguous quote given the event context and author’s affiliation. E.g., “protect our children from mass shootings” could easily be disambiguated as either “ban guns” or “arm teachers” when the author’s stance on the issue of ‘gun rights’ is known.
Our data collection effort is designed to capture different aspects of social context grounding and facilitate detailed error analysis. Defined as a binary classification task over tuples ⟨Party, Event, Vague text, Explicit text⟩, we create negative examples by flipping tuple elements values of positive examples. This allows us to evaluate whether models can capture event relevance, political stance, or constrain the interpretation based on the vague text. For example, in the context of Event #1 in fig. 1, we can test if models simply capture the correlation between Democrats and negative stance towards guns access by replacing the vague text to “let your voice be heard”, which would make the interpretation in fig. 1 implausible despite being consistent with that stance, while other consistent interpretations would be plausible (e.g., “go outside and join the march for our lives”).
3.2.1 Vague Text Data Collection
Data collection was done in several steps. (1)Vague Texts Collection. We collected vague text candidates from tweets by US politicians (i.e. senators and representatives) between the years $20192019$ to $20212021$ from Congress Tweets corpus. We identified a list of $99$ well-known events from that period and identified event-related tweets using their time frame and relevant hashtags. We used a pre-trained BERT-based (Devlin et al., 2019) NER model to collect tweets that do not contain any entity mentions to identify potential candidates for vague texts. We manually identified examples that could have contrasting senses by flipping their social context. We obtain $9393$ vague text candidates via this process.
(2)In-Context Plausible Meaning Annotation. We match the $9393$ ambiguous tweets with different events that fit them. We use both Democrat and Republican as the author party affiliation. We obtain $600600$ context-tweet pairs for AMT annotation. For each tweet, we ask AMT workers to annotate the following two aspects: 1) sentiment towards the three most relevant entities in the event (sanity check) and 2) a detailed explanation of the intended meaning given the event and author’s party affiliation. We obtain $469469$ reasonable annotations. After this step, each annotation was screened by in-house annotators. We ask three in-house annotators to vote on the correctness, appropriateness, and plausibility of the annotation given the context. Thus, we create a total of $374374$ examples.
(3)LLM-based Data Expansion. Using these examples, we further generate candidates for the task using LLM few-shot prompting. We use the examples from the previous step as in-context few-shot examples in the prompt. We use GPT-NeoX (Black et al., 2022) and GPT-3 (Brown et al., 2020a) for candidate generation. For each generated answer, manual inspection by three in-house annotators is performed to ensure data quality. We generate $928928$ candidates using GPT-NeoX and GPT-3. Human expert filtering results in $650650$ generations that pass the quality check. After removing redundant samples, we obtain $365365$ examples. Thus, we obtain a total of $739739$ annotations for this task. Then, for each of the $739739$ examples, we ask in-house annotators to select $33$ relevant negative options from the pool of explanations. We instruct them to pick hard examples that might contain similar entities as the correct interpretation. This results in 2,95629562,9562 , 956 binary classification data samples. We analyze and discuss the results of human validation of large LM generations in section 6.
Similar to the previous task, we split the train, test sets by events, and vague text to test the general social understanding capabilities of the model. We reserve Donald Trump’s second impeachment verdict event for the test set. We also reserve Democratic examples of $22$ events and Republican examples of $22$ events exclusively for the test set. We split the dataset into 1,91619161,9161 , 916 train, $460460$ development, and $580580$ test examples. $180180$ of the test examples are from events/party contexts unseen in train data.
4 Modeling Social Context
Model Target Identification Sentiment Identification
Prec Rec Macro-F1 Acc Prec Rec Macro-F1 Acc
No Context Baselines BERT-large 69.09 72.35 68.83 70.56 58.74 60.17 58.95 58.37
RoBERTa-base 66.58 69.54 65.14 66.40 61.68 61.27 61.36 60.65
PLMs +Twitter Bio Context BERT-large + user-bio 69.03 71.86 69.34 71.66 60.02 60.44 60.13 59.86
RoBERTa-base + user-bio 65.83 68.65 64.79 66.30 60.06 59.91 59.94 59.46
PLMs +Wikipedia Context BERT-large + wiki 63.58 65.78 60.33 61.05 53.48 56.44 53.9 53.32
RoBERTa-base + wiki 69.02 72.32 68.62 70.27 57.62 59.10 58.07 58.28
LLMs GPT-3 0-shot 69.25 70.58 69.77 73.78 56.20 55.04 54.18 56.80
GPT-3 4-shot 69.81 72.99 66.45 67.03 58.12 57.10 55.00 57.51
Static Contextutalized Embedding Models RoBERTa-base + PAR Embs 68.38 71.63 67.67 69.18 55.01 56.89 55.51 55.40
BERT-large + PAR Embs 65.40 67.33 60.25 60.56 55.24 57.54 55.89 55.80
RoBERTa-base + DCF Embs 72.89 75.95 73.56 75.82 63.05 63.52 62.90 63.03
BERT-large + DCF Embs 68.76 72.02 68.32 69.97 61.59 63.25 61.22 60.75
Discourse Contextualized Models BERT-large + DCF 71.12 74.61 71.17 72.94 65.81 65.25 65.34 65.31
RoBERTa-base + DCF 70.44 73.86 70.39 72.15 63.45 63.34 63.37 63.23
Table 2: Results of baseline experiments on Target Entity (binary task) and Sentiment ($44$-classes) test sets. We report macro-averaged Precision, macro-averaged Recall, macro-averaged F1, and Accuracy metrics.
Model Vague Text Disambiguation
Prec Rec Macro-F1 Acc
No Context Baselines
BERT-large $52.2452.24$ $55.5855.58$ $50.2850.28$ $53.7553.75$
RoBERTa-base $55.355.3$ $51.8251.82$ $54.5354.53$ $56.0856.08$
PLMs + Wikipedia Context
BERT-large + wiki $52.3152.31$ $46.9046.90$ $66.8766.87$ $76.0376.03$
BERT-base + wiki $51.8551.85$ $38.6238.62$ $64.3664.36$ $75.6975.69$
LLMs
GPT-3 0-shot 63.10 $62.9262.92$ $62.5862.58$ $63.563.5$
GPT-3 4-shot $62.0562.05$ $62.2962.29$ $61.8661.86$ $62.0462.04$
Static Contextutalized Embedding Models
BERT-large + PAR $47.6847.68$ $49.6649.66$ $65.5365.53$ $73.7973.79$
BERT-base + PAR $45.9345.93$ $54.4854.48$ $65.4965.49$ $72.5972.59$
BERT-large + DCF Embs $47.1847.18$ 63.45 $67.5567.55$ $73.1073.10$
BERT-base + DCF Embs $56.5856.58$ $59.3159.31$ 71.71 78.45
Discourse Contextualization Models
BERT-large + DCF $52.7652.76$ $59.3159.31$ $69.9469.94$ $76.5576.55$
BERT-base + DCF $52.7352.73$ $60.0060.00$ $70.0670.06$ $76.5576.55$
Table 3: Results of baseline experiments on Vague Text Disambiguation dataset test split, a binary classification task. We report macro-averaged Precision, macro-averaged Recall, macro-averaged F1, and Acc. metrics
The key technical question this paper puts forward is how to model the social context, such that the above tasks can be solved with high accuracy. We observe that humans can perform this task well (section 6.3), and evaluate different context modeling approaches in terms of their ability to replicate human judgments. These correspond to No Context, Text-based context representation (e.g., Twitter Bio, relevant Wikipedia articles), and Graph-based context representation, simulating the social media information that human users are exposed to when reading the vague texts.
We report the results of all our baseline experiments in table 2 and table 3. The first set of results evaluate fine-tuned pre-trained language models (PLM), namely BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019), with three stages of modeling context. Firstly, we evaluate no contextual information setting. Second, we include the authors’ Twitter bios as context. Finally, we evaluate the information from the author, event, and target entity Wikipedia pages as context (models denoted PLM Baselines {No, Twitter Bio, Wikipedia} Context, respectively).
We evaluate GPT-3333gpt-3.5-turbo-1106 via OpenAI API in zero-shot and four-shot in-context learning paradigm on both tasks. We provide contextual information in the prompt in the form of short event descriptions and authors’ affiliation descriptions. Note that GPT-3 is trained on news data until Sep. $20212021$ which includes the events in our data (models denoted LLM Baseline).
We evaluate the performance of politician embeddings from Political Actor Representation (PAR) (Feng et al., 2022) and Discourse Contextualization Framework (DCF) (Pujari and Goldwasser, 2021) models. (models denoted Static Contextutalized Embeddings). We use PAR embeddings available on their GitHub repository [^4] https://github.com/BunsenFeng/PAR. For DCF model, we use released pre-trained models from GitHub repository [^5] https://github.com/pujari-rajkumar/compositional_learner to generate author, event, text, and target entity embeddings. We evaluate the embeddings on both tasks. We briefly review these models in section 4.1 & section 4.2.
Finally, we use tweets of politicians from related previous events and build context graphs for each data example as proposed in Pujari and Goldwasser (2021). We use Wikipedia pages of authors, events, and target entities to add social context information to the graph. Then, we train the Discourse Contextualization Framework (DCF) for each task and evaluate its performance on both tasks (models denoted Discourse Contextualization Model). Further details of our baseline experiments are presented in subsection section 4.3. Results of our baseline experiments are discussed in section 5.
4.1 Discourse Contextualization Framework
Discourse Contextualization Framework (DCF) (Pujari and Goldwasser, 2021) leverages relations among social context components to learn contextualized representations for text, politicians, events, and issues. It consists of encoder and composer modules that compute holistic representations of the context graph. The encoder creates an initial representation of nodes. Composer propagates the information within the graph to update node representations. They further define link prediction learning tasks over context graphs to train the model. They show that their representations significantly outperform several PLM-based baselines trained using the same learning tasks. Their qualitative evaluation shows that these representations capture interesting aspects of context information.
4.2 Political Actor Representation
Feng et al. (2022) propose the Political Actor Representation (PAR) framework, a graph-based approach to learn more effective politician embeddings. They propose three learning tasks, namely, $11$) Expert Knowledge Alignment $22$) Stance Consistency training & $33$) Echo chamber simulation, to infuse social context into the politician representations. They show that PAR representations outperform state-of-the-art models on Roll Call Vote Prediction and Political Perspective Detection.
4.3 Experimental Setup
Target Entity Detection is binary classification with ⟨author, event, tweet, target-entity⟩as input and target/non-target label as output. Sentiment Detection is set up as $44$-way classification. Input is the same as the target task and output is one of: {positive, neutral, negative, non-target}. Vague Text Disambiguation is a binary classification task with ⟨party-affiliation, event, vague-text, explanation-text⟩and a match/no-match label as output.
In phase 1 no-context baselines, we use the author, event, tweet, and target embeddings generated by PLMs. We concatenate them for input. In Twitter-bio models, we use the author’s Twitter bio embeddings to represent them. Wiki context models receive Wikipedia page embeddings of author, event, and target embeddings. It is interesting to note that the Wikipedia context models get all the information needed to solve the tasks.. In phase 2 LLM experiments, we use train samples as in-context demonstrations. We provide task and event descriptions in the prompt. In phase 3 PAR models, we use politician embeddings released on the PAR GitHub repository to represent authors. We replace missing authors with their wiki embeddings. For the Vague Text task, we average PAR embeddings for all politicians of the party to obtain party embeddings. For DCF embedding models, we generate representations for all the inputs using context graphs. We also use authors’ tweets from relevant past events. We build graphs using author, event, tweet, relevant tweets, and target entity as nodes and edges as defined in the original DCF paper. In phase 4, we use the same setup as the DCF embedding model and additionally back-propagate to DCF parameters. This allows us to fine-tune the DCF context graph representation for our tasks.
5 Results
We show the results of our baseline experiments in tables 2 and 3. We evaluate our models using macro-averaged precision, recall, F1, and accuracy metrics. As the classes are not balanced, macro-F1 is the main metric we observe. Overall, several patterns emerge from the results consistent across all tasks. First, modeling social context is still an open problem. None of our models were able to perform close to human level. Second, adding context can help performance, compared to the No-Context baselines, models incorporating context performed better, with very few exceptions. Third, LLMs are not the panacea for social-context pragmatic tasks. Despite having access to a textual context representation as part of the prompt, and having access to relevant event-related documents during their training phase, these models under-perform compared to much simpler models that were fine-tuned for this task. Finally, explicit context modeling using the DCF model consistently leads to the best performance. Across all of our tasks, some variation of DCF-based model provided the best performance. The DCF model mainly represents the social context in the form of text documents for all nodes. Further symbolic addition of other types of context such as social relationships among politicians and relationships between various nodes could further help in achieving better performance on these tasks. In the Target Entity task, RoBERTa-base + DCF embeddings obtain $73.5673.56$ F1 vs. $68.8368.83$ for the best no-context baseline. Twitter bio and wiki-context hardly improve, demonstrating the effectiveness of modeling contextual information explicitly vs. concatenating context as text documents. No context performance well above the random performance of $5050$ F1 indicates the bias in the target entity distribution among classes. We discuss this in section 6.4. In Sentiment Identification task, we see that BERT-large + DCF back-propagation outperforms all the other models. Vague Text Disambiguation task results in table 3 show that DCF models outperform other models significantly. $71.7171.71$ F1 is obtained by BERT-base + DCF embeddings. BERT-base performing better than bigger PLMs might be due to DCF model’s learning tasks being trained using BERT-base embeddings.
6 Analysis and Discussion
Democrat Only Entities Common Entities Republican Only Entities
Target Sentiment Agreed-Upon Entities Divisive Entities Target Sentiment
Target Sentiment Sentiment (D) Target Sentiment (R)
Anita Hill
Patty Murray
Merrick Garland
Jeff Flake
Positive
Positive
Positive
Negative
US Supreme Court
US Senate
FBI
Judiciary Committee
Neutral
Neutral
Neutral
Neutral
Positive
Positive
Positive
Negative
Negative
Negative
Christine Blasey Ford
Deborah Ramirez
Julie Swetnick
Brett Kavanaugh
Donald Trump
Mitch McConnell
Negative
Negative
Negative
Positive
Positive
Positive
Susan Collins
Chuck Grassley
Diane Feinstein
Chuck Schumer
Sean Hannity
Positive
Positive
Negative
Negative
Neutral
Table 4: Target Entity-Sentiment centric view of Kavanaugh Supreme Court Nomination discourse
6.1 Ablation Analysis on Vague Text Task
We report ablation studies in table 5 on the Vague Text task test set. We consider $55$ splits: (1) Unseen Party: ⟨party, event⟩not in the train set but ⟨opposing-party, event⟩is present, (2) Unseen Event: ⟨party ⟩not in train set, (3) Flip Event: negative samples with corresponding ‘event flipped-party/vague tweet matched’ positive samples in train set and analogous (4) Flip Party and (5) Flip Tweet splits. We report the representatives from each category. We observe that all models obtain weaker performance on unseen splits, as expected, unseen events being the hardest. Contextualized models achieve higher margins. DCF gains 7.6(13.2%) and DCF embeddings attain 8.12(20.42%) macro-F1 improvement over BERT-base+wiki compared to respective margins of 8.86% and 11.42% on the full test set. In the flip splits with only negative examples, accuracy gain over random baseline for all splits is seen. This indicates that models learn to jointly condition on context information rather than learn spurious correlations over particular aspects of the context. Specifically, flip-tweet split results indicate that models don’t just learn party-explanation mapping.
Data Split
Unseen
Party
Unseen
Event
Flip
Tweet
Flip
Event
Flip
Party
Ma-F1 Ma-F1 Acc Acc Acc
Random 44.70 29.69 75 75 75
BERT-base+wiki 57.58 39.76 88.14 89.77 87.77
BERT-base
+DCF Embs
61.79 47.88 86.10 93.18 84.57
BERT-base+DCF 65.18 45.65 82.03 89.77 84.04
Table 5: Ablation Study Results on Vague Text Task
6.2 Vague Text LLM Generation Quality
[…]
6.3 Vague Text Human Performance
[…]
6.4 Target Entity Visualization
In table 4, we study an event, Kavanaugh Supreme Court Nomination, using entity target sentiment annotations. We identify entities that are discussed by both parties. We further separate them into divisive and agreed-upon entities based on expressed sentiments. We also show partisan discussed entities for the event. This analysis paints a very accurate picture of the discussed event. We observe that the main entities of Trump, Dr. Ford, Kavanaugh, Sen. McConnell, and other accusers/survivors emerge as divisive entities. Entities such as Susan Collins and Anita Hill who were vocal mouthpieces of the respective party stances but didn’t directly participate in the event emerge as partisan entities. Supreme Court, FBI, and other entities occur in the discourse but only as neutral entities. This analysis shows the usefulness of the Target Entity and Sentiment identification in accurately summarizing events. Successful models on these tasks could greatly help in understanding political events better.
6.5 DCF Context Understanding
We look into examples that are incorrectly predicted using Wikipedia pages but correctly predicted when using the DCF model. We report some of the examples in table 6 in the appendix. In examples $11$ & $22$ of Target Entity-Sentiment task, we see that when the entity is not explicitly mentioned in the tweet, the Wiki-Context model fails to identify them as the target entities. We posit that while the Wikipedia page of each relevant event will contain these names, explicit modeling of entities in DCF model allows these examples to be correctly classified. Examples 13131-31 - 3 of Vague Text Disambiguation task show that when no clear terms indicate the sentiment towards a view, the Wiki-Context model fails to disambiguate the tweet text. Explicit modeling of politician nodes in the DCF model might allow these texts to be correctly disambiguated.
7 Conclusion and Future Work
In this paper, we motivate, define, and operationalize “Social Context Grounding” for political discourse. We build two datasets that are useful to evaluate social context grounding in NLP models. We experiment with many types of context modeling and benchmark SOTA models. We show that explicit modeling of social context outperforms naive concatenation of context and LLMs such as GPT-3. Future work includes building datasets for other components of Social Context Grounding and models that account for context in various forms.
Limitations
Our work only addresses English language text in US political domain. We also build upon large language models and large PLMs which are trained upon huge amounts of uncurated data. Although we employed human validation at each stage, biases could creep into the datasets. We also don’t account for the completeness of out datasets as it is a pioneering work on a new problem. Social context is vast and could have myriad of components. We only take a step in the direction of social context grounding in this work. The performance on these datasets might not indicate full social context understanding but they should help in sparking research in the direction of models that explicitly model such context. Although we tuned our prompts a lot, it is possible that better prompts and evolving models might produce better results on the LLM baselines. Our qualitative analysis is predicated on a handful of examples. They are attempts to interpret the results of large neural models and hence don’t carry as much confidence as our empirical observations.
Ethics Statement
In this work, our data collection process consists of using both AMT and GPT-3. For the Target Entity and Sentiment task, we pay AMT workers \$1currency-dollar1\\$1\$ 1 per HIT and expect an average work time of $33$ minutes. This translates to an hourly rate of \$20currency-dollar20\\$20\$ 20 which is above the federal minimum wage. For the Vague Text Disambiguation task, we pay AMT workers \$1.10currency-dollar1.10\\$1.10\$ 1.10 per HIT and expect an average work time of $33$ minutes. This translated to an hourly rate of \$22currency-dollar22\\$22\$ 22.
We recognize collecting political views from AMT and GPT-3 may come with bias or explicit results and employ expert gatekeepers to filter out unqualified workers and remove explicit results from the dataset. Domain experts used for annotation are chosen to ensure that they are fully familiar with the events in focus. Domain experts were provided with the context related to the events via their Wikipedia pages, background on the general issue in focus, fully contextualized quotes, and authors’ historical discourse obtained from ontheissues.org. We have an annotation quid-pro-quo system in our lab which allows us to have a network of in-house annotators. In-house domain experts are researchers in the CSS area with familiarity with a range of issues and stances in the US political scene. They are given the information necessary to understand the events in focus in the form of Wikipedia articles, quotes from the politicians in focus obtained from ontheissues.org, and news articles related to the event. We make the annotation process as unambiguous as possible. In our annotation exercise, we ask the annotators to mark only high-confidence annotations that can be clearly explained. We use a majority vote from $33$ annotators to validate the annotations for the target entity task.
Our task is aimed at understanding and grounding polarized text in its intended meaning. We take examples where the intended meaning is clearly backed by several existing real-world quotes. We do not manufacture the meaning to the vague statements, we only write down unambiguous explanations where context clearly dictates the provided meaning. Applications of our research as we envision would be adding necessary context to short texts by being able to identify past discourse from the authors that are relevant to the particular text in its context. It would also be able to ground the text in news articles that expand upon the short texts to provide full context.
Appendix A Reproducibility
We use the HuggingFace Transformers (Wolf et al., 2020) library for PLMs. We use GPT-NeoX implementation by ElutherAI (Black et al., 2022) and GPT-3 (Brown et al., 2020b) via OpenAI API for our LLM baselines. We run $100100$ epochs for all experiments. We use $1010$ NVIDIA GeForce 1080i GPUs for our experiments. We use the train, development, and test splits detailed in section 3 for our experiments. We use the development macro-F1 for early stopping. We run all our experiments using random seeds to ensure reproducibility. We experiment with a random seed value set to {13131313}. We report the results of the $33$ fold cross-validation. We report only the mean on all the cross-fold validation results for clarity. All our code, datasets, and result logs will be released publicly upon acceptance. We experiment with $33$, $55$, and $1010$ fold cross-validation. As the results on the development data are almost identical, we report the results of $33$ fold cross-validation in all our experiments.
Appendix B Error Analysis
Target Entity and Sentiment Task Vague Text Disambiguation Task
Tweet: Republicans held Justice Scalia’s seat open for more
than 400 days. Justice Kennedy’s seat has been vacant for
less than two months. It’s more important to investigate a
serious allegation of sexual assault than to rush Kavanaugh
onto the Supreme Court for a lifetime appointment.
Tweet: Thanks for this.
Author: Adam Schiff (Democrat) Affiliation: Democrat
Event: Brett Kavanaugh Supreme Court nomination Event: United States withdrawal from the Paris Agreement
Entity: Christine Blasey Ford
Paraphrase: There’s nothing surprising in withdrawing from
the Paris agreement. Thanks for not caring our environment and
future generations.
Wiki-Context Prediction: Not Target | DCF Prediction: Target (correct) Wiki-Context Prediction: No | DCF Prediction: Yes (correct)
Tweet: We will not be intimidated. Democracy will not be
intimidated. We must hold the individuals responsible for the
Jan. 6th attack on the U.S. Capitol responsible. Thank you
@RepAOC for tonight’s Special Order Hour and we will
continue our efforts to #HoldThemAllAccountable.
Tweet: Let us say enough. Enough.
Author: Adriano Espaillat (Democrat) Affiliation: Democrat
Event: January 6 United States Capitol attack
Event: Second impeachment of Donald Trump ended with not
guilty
Entity: Donald Trump
Paraphrase: The failure of the Democrats to impeach Donald
Trump is a strong moment for our legislature which can get
back to its work helping the American people. Today we’ve been
able to tell the American people what we have known all along,
that Donald Trump was not guilty of these charges.
Wiki-Context Predicted: Not Target | DCF Prediction: Target (correct) Wiki-Context Predicted: Yes | DCF Prediction: No (correct)
Tweet: #GeorgeFloyd #BlackLivesMatter #justiceinpolicing
QT @OmarJimenez Former Minneapolis police officer Derek
Chauvin is in the process of being released from the Hennepin
County correctional facility his attorney tells us. He is one of
the four officers charged in the death of George Floyd. He
faces murder and manslaughter charges.
Tweet: Lots of honking and screaming from balconies.
Something must be going on.
Author: Adriano Espaillat (Democrat) Affiliation: Democrat
Event: George Floyd protests Event: Presidential election of 2020
Entity: Derek Chauvin
Paraphrase: I’m sure that the people are celebrating the
election results.
Wiki-Context Predicted Sentiment: Positive | DCF Prediction: Negative (correct) Wiki-Context Prediction: No | DCF Prediction: Yes (correct)
Table 6: Examples where baseline model fails but DCF works
Appendix C Annotation Interfaces
Figure 2: An example of Tweet Target Entity and Sentiment Annotation GUI
Figure 3: An example of Vague Text Disambiguation GUI

### v3

@@
 Abstract
-Social media discourse frequently consists of ‘seemingly similar language used by opposing sides of the political spectrum’, often translating to starkly contrasting perspectives.
-E.g., “thoughts and prayers”, could express sympathy for mass-shooting victims, or criticize the lack of legislative action on the issue.
-This paper defines the context required to fully understand such ambiguous statements in a computational setting and ground them in real-world entities, actions, and attitudes.
-We propose two challenging datasets that require an understanding of the real-world context of the text.
-We benchmark these datasets against models built upon large pre-trained models, such as RoBERTa and GPT-3.
-Additionally, we develop and benchmark more structured models building upon existing Discourse Contextualization Framework and Political Actor Representation models.
-We analyze the datasets and the predictions to obtain further insights into the pragmatic language understanding challenges posed by the proposed social grounding tasks.
-“We Demand Justice!”: Towards Social Context Grounding of Political Texts
-Rajkumar Pujari Purdue University, USA rpujari@purdue.edu                        Chengfei Wu Purdue University, USA wu1491@purdue.edu                        Dan Goldwasser Purdue University, USA dgoldwas@purdue.edu
+Political discourse on social media often contains similar language with opposing intended meanings.
+For example, the phrase thoughts and prayers, is used to express sympathy for mass shooting victims, as well as satirically criticize the lack of legislative action on gun control.
+Understanding such discourse fully by reading only the text is difficult.
+However, knowledge of the social context information makes it easier.
+We characterize the social context required to fully understand such ambiguous discourse, by grounding the text in real-world entities, actions, and attitudes.
+We propose two datasets that require an understanding of social context and benchmark them using large pre-trained language models and several novel structured models.
+We show that structured models, explicitly modeling social context, outperform larger models on both tasks, but still lag significantly behind human performance.
+Finally, we perform an extensive analysis, to obtain further insights into the language understanding challenges posed by our social grounding tasks.
+“We Demand Justice!”:
+Towards Social Context Grounding of Political Texts
+Rajkumar Pujari  and Chengfei Wu  and Dan Goldwasser Purdue University, West Lafayette, USA {rpujari,wu1491,dgoldwas}@purdue.edu
 1 Introduction
-Over the past decade micro-blogging websites have become the primary channel of engagement between politicians and people in the US, resulting in an increasingly succinct political discourse.
-Politicians from the same party often coordinate the phrasing of social messaging in developing events to amplify their impact (Vaes et al., 2011; Weber and Neumann, 2021).
-Hence, repetitive phrases such as ‘Thoughts and Prayers’ are extensively used to signal more nuanced stances across several events.
-The interaction among politicians from opposing parties often leads to messaging phrased similarly but intended to signal opposing real-world actions.
-‘Thoughts and Prayers’ is used to express condolences by Republicans in mass shooting events, and by Democrats convey an angry or sarcastic tone as a call for action demanding ‘tighter gun control measures’.
-The example in fig.
+Over the past decade, micro-blogging websites have become the primary medium for US politicians to interact with general citizens and influence their stances for gaining support.
+As a result, politicians from the same party often coordinate the phrasing of their social messaging, to amplify their impact (Vaes et al., 2011; Weber and Neumann, 2021).
+Hence, repetitive, succinct phrases, such as “Thoughts and Prayers”, are extensively used, although they signal more nuanced stances.
+Moreover, the interaction among politicians from opposing parties often leads to messaging phrased similarly, but signaling opposing real-world actions.
+For example, ‘Thoughts and Prayers’, when used by Republicans, expresses condolences in mass shooting events, but when used by Democrats conveys an angry or sarcastic tone as a call for action demanding “tighter gun control measures”.
+Similarly, fig.
 1 shows contrasting interpretations of the phrase “We need to keep our teachers safe!” depending on different speakers and in the context of different events.
@@
 Table 1: Examples of Annotated Datasets and their statistics
-Humans familiar with the stances of a politician and, possessing knowledge about the event from the news, can easily understand the intended meaning.
-But, computationally understanding the correct meaning of such language is challenging.
+Humans familiar with the stances of a politician and, possessing knowledge about the event from the news, can easily understand the intended meaning of political phrases.
+However, automatically understanding such language is challenging.
 Our main question in this paper is - Can an NLP model find the right meaning?
-From a linguistic perspective, we follow the distinction (Bach, 2008) between semantic interpretation (where the meaning is encoded in the utterance directly and does not change based on its external context), and pragmatic interpretation (that depends on extra-linguistic information).
+From a linguistic perspective, we follow the distinction (Bach, 2008) between semantic interpretation (i.e., meaning encoded directly in the utterance and does not change based on its external context), and pragmatic interpretation (that depends on extra-linguistic information).
 The latter has gathered significant interest in the NLP community recently (Bender and Koller, 2020; Bisk et al., 2020), focusing on language understanding, when grounded in an external context (Fried et al., 2023).
-To a large extent, the focus of such studies was on grounding language in a perceptual environment (e.g., image captioning (Andreas and Klein, 2016; Sharma et al., 2018; Alikhani et al., 2020), instruction following (Wang et al., 2016; Suhr et al., 2019; Lachmy et al., 2022), and game playing (Potts, 2012; Udagawa and Aizawa, 2019) tasks).
-Unlike these works, in this paper, we focus on grounding language in a social context, i.e., modeling the common ground (Clark and Brennan, 1991; Traum, 1994; Stalnaker, 2002) between the author and their social media followers that enables understanding what would otherwise be highly ambiguous utterances.
+To a large extent, the focus of such studies has been on grounding language in a perceptual environment (e.g., image captioning (Andreas and Klein, 2016; Sharma et al., 2018; Alikhani et al., 2020), instruction following (Wang et al., 2016; Suhr et al., 2019; Lachmy et al., 2022), and game playing (Potts, 2012; Udagawa and Aizawa, 2019) tasks).
+Unlike these works, in this paper, we focus on grounding language in a social context, i.e., modeling the common ground (Clark and Brennan, 1991; Traum, 1994; Stalnaker, 2002) between the author and their social media followers, that enables understanding an otherwise highly ambiguous utterances.
 The Social Context Understanding, needed for building successful models for such tasks, can come from a wide variety of sources.
-The politician’s party affiliation, historical stances on the issue, social relationships, knowledge about the involved entities, and related prior and upcoming events, are a few examples of the relevant social context.
+The politician’s affiliation and historical stances on the issue provide can capture crucial social context.
+Social relationships, knowledge about the involved entities, and related prior and upcoming events form important part of the puzzle as well.
 In fig.
-1 event #1#1\#1# 1, combining the event information (school shooting) with the speakers’ stances on gun control, would facilitate understanding the intended meaning of the text better.
-This paper operationalizes the problem of ‘Social Context Grounding’ by defining tasks capturing the intended meaning at different level of granularity.
-At the most basic level, providing the social context should help identify the entities targeted and the sentiment towards them.
+1 event #1#1\#1# 1, combining the event information (school shooting) with the speakers’ gun control stances, would facilitate understanding the intended meaning of the text.
+The main motivation of this paper work is to operationalize the ‘Social Context Grounding’ problem as a pragmatic understanding task.
+From a practical perspective, this would enable the creation of better NLP-CSS models that can process social media text in settings that require contextualized understanding.
+We suggest several datasets, designed to evaluate this ability in computational models.
+These task capture the intended meaning at different level of granularity.
+At the most basic level, providing the social context can help identify the entities targeted, and the sentiment towards them.
 In fig.
-1, the social context ⟨Ev.#1, Harris⟩and the text “we need to keep our teachers safe” $\Rightarrow$ negative attitude towards guns.
+1, the social context ⟨event#1, Harris⟩and the text “we need to keep our teachers safe” $\Rightarrow$ negative attitude towards guns.
 A more nuanced account of meaning, which we formulate as a separate task, captures the specific means in which the negative attitude is expressed (the Interpretation in fig.
 1).
-We present two datasets corresponding to these tasks, namely, ‘Target Entity and Sentiment Detection’ and ‘Vague Text Disambiguation’.
+We additionally present two datasets corresponding to these tasks, namely, ‘Target Entity and Sentiment Detection’ and ‘Vague Text Disambiguation’.
 In the first, the goal is to predict: $11$) whether a given entity is the intended target of a politician’s tweet and $22$) the sentiment towards the intended targets.
@@
 These include pre-trained LM-based classifiers, and LLM in-context learning (Brown et al., 2020a; Black et al., 2022), which use a textual representation of the social context.
-We adopt an existing graph-based discourse contextualization framework (Pujari and Goldwasser, 2021; Feng et al., 2022) to explicitly model the social context needed to solve the proposed tasks.
-Our results demonstrate that discourse contextualization models outperform other models on both tasks.
-We present an error analysis of the outputs to gain further insights.
+We also adopt an existing graph-based discourse contextualization framework (Pujari and Goldwasser, 2021; Feng et al., 2022), to explicitly model the social context needed to solve the proposed tasks.
+Our results demonstrate that the discourse contextualization models outperform other models on both tasks.
+We present an error analysis to gain further insights.
 We describe the models in section 4 and the results in section 5.
-We also present qualitative visualization of a political event, Brett Kavanaugh Supreme Court Nomination in section 6.4 from the lens of target entity-sentiment to showcase a unique summary of the event discourse.
-Then, we perform human evaluation on ‘Vague Text Disambiguation’ and observe that humans find the task much easier than the evaluated models.
+We also present a qualitative visualization of a political event, Brett Kavanaugh Supreme Court Nomination (section 6.4), from target entity-sentiment perspective.
+It showcases a unique summary of the event discourse.
+We perform human evaluation on our ‘Vague Text Disambiguation’ dataset, and observe that humans find this task much easier than the evaluated models.
 We also present observations of human vs.
-LLM errors in disambiguations.
-Our contributions can be summarized as:
+LLM errors in disambiguation.
+In summary, our contributions are:
   1. 1.
@@
     Evaluating various state-of-the-art context representation models on the task.
-Adopting existing discourse contextualization framework for the proposed tasks and evaluating GPT-3 in-context learning performance on the tasks
+We adopt existing discourse contextualization framework for the proposed tasks, and evaluate GPT-3’s in-context learning performance, as well.
   3. 3.
-    Performing human studies to benchmark the dataset difficulty and GPT-3 generation performance comparison with human workers [^1] Our data and code will be released under MIT license
+    Performing human studies to benchmark the dataset difficulty and GPT-3 generation performance, when compared to human workers.
+[^1] Our data and code is at https://github.com/pujari-rajkumar/language-in-context
 2 Related Work
@@
 We split the examples into 4,37043704,3704 , 370 train, $511511$ development, and 1,00910091,0091 , 009 test examples.
+We compute the mean Cohen’s kappa score for annotations and report inter-annotator agreement for annotated targets (0.47) and sentiment (0.73)
 3.2 Vague Text Disambiguation Task
@@
 We identified a list of $99$ well-known events from that period and identified event-related tweets using their time frame and relevant hashtags.
-We used a pre-trained BERT-based (Devlin et al., 2019) NER model to collect tweets that do not contain any entity mentions to identify potential candidates for vague texts.
+We used a pre-trained BERT-based (Devlin et al., 2019) NER model to collect tweets that contain few or no entity mentions to identify potential candidates for vague texts.
 We manually identified examples that could have contrasting senses by flipping their social context.
@@
 We match the $9393$ ambiguous tweets with different events that fit them.
-We use both Democrat and Republican as the author party affiliation.
+We use both Democrat/Republican as the author party affiliation.
 We obtain $600600$ context-tweet pairs for AMT annotation.
@@
 We use GPT-NeoX (Black et al., 2022) and GPT-3 (Brown et al., 2020a) for candidate generation.
-For each generated answer, manual inspection by three in-house annotators is performed to ensure data quality.
+Manual inspection by three in-house annotators is performed for each generated answer to ensure data quality.
 We generate $928928$ candidates using GPT-NeoX and GPT-3.
-Human expert filtering results in $650650$ generations that pass the quality check.
-After removing redundant samples, we obtain $365365$ examples.
+Manual filtering results in $650650$ generations that pass the quality check.
+After removing redundant samples, we obtain $365365$ additional examples.
 Thus, we obtain a total of $739739$ annotations for this task.
 Then, for each of the $739739$ examples, we ask in-house annotators to select $33$ relevant negative options from the pool of explanations.
-We instruct them to pick hard examples that might contain similar entities as the correct interpretation.
+We instruct them to pick hard examples that potentially contain overlapping entities with the gold answer.
 This results in 2,95629562,9562 , 956 binary classification data samples.
-We analyze and discuss the results of human validation of large LM generations in section 6.
+We analyze and discuss the results of human validation of large LM generations in section 6).
+This process allows us to create three variants of the task: binary-classification, multiple-choice and generation variants.
+We evaluate several classification models on the binary classification variant (Tab.3).
+We evaluate LLMs on the generation variant (§6.2).
+We benchmark humans and the best models on the multiple-choice variant (§6.3).
 Similar to the previous task, we split the train, test sets by events, and vague text to test the general social understanding capabilities of the model.
@@
 We evaluate GPT-3333gpt-3.5-turbo-1106 via OpenAI API in zero-shot and four-shot in-context learning paradigm on both tasks.
-We provide contextual information in the prompt in the form of short event descriptions and authors’ affiliation descriptions.
+We provide contextual information in the prompt as short event descriptions and authors’ affiliation descriptions.
 Note that GPT-3 is trained on news data until Sep.
@@
 Composer propagates the information within the graph to update node representations.
-They further define link prediction learning tasks over context graphs to train the model.
+They define link prediction learning tasks over context graphs to train the model.
 They show that their representations significantly outperform several PLM-based baselines trained using the same learning tasks.
-Their qualitative evaluation shows that these representations capture interesting aspects of context information.
 4.2 Political Actor Representation
@@
 They propose three learning tasks, namely, $11$) Expert Knowledge Alignment $22$) Stance Consistency training & $33$) Echo chamber simulation, to infuse social context into the politician representations.
-They show that PAR representations outperform state-of-the-art models on Roll Call Vote Prediction and Political Perspective Detection.
+They show that PAR representations outperform SOTA models on Roll Call Vote Prediction and Political Perspective Detection.
 4.3 Experimental Setup
@@
 5 Results
-We show the results of our baseline experiments in tables 2 and 3.
-We evaluate our models using macro-averaged precision, recall, F1, and accuracy metrics.
-As the classes are not balanced, macro-F1 is the main metric we observe.
-Overall, several patterns emerge from the results consistent across all tasks.
+The results of our baseline experiments are described in Tab.
+2 and 3.
+We evaluate our models using macro-averaged precision, recall, F1, and accuracy metrics (due to class imbalance, we focus on macro-F1).
+Several patterns, consistent across all tasks, emerge.
 First, modeling social context is still an open problem.
@@
 Finally, explicit context modeling using the DCF model consistently leads to the best performance.
-Across all of our tasks, some variation of DCF-based model provided the best performance.
 The DCF model mainly represents the social context in the form of text documents for all nodes.
@@
 We discuss this in section 6.4.
-In Sentiment Identification task, we see that BERT-large + DCF back-propagation outperforms all the other models.
+In Sentiment Identification task, we see that BERT-large + DCF back-propagation outperforms all other models.
 Vague Text Disambiguation task results in table 3 show that DCF models outperform other models significantly.
@@
 We consider $55$ splits: (1) Unseen Party: ⟨party, event⟩not in the train set but ⟨opposing-party, event⟩is present, (2) Unseen Event: ⟨party ⟩not in train set, (3) Flip Event: negative samples with corresponding ‘event flipped-party/vague tweet matched’ positive samples in train set and analogous (4) Flip Party and (5) Flip Tweet splits.
-We report the representatives from each category.
-We observe that all models obtain weaker performance on unseen splits, as expected, unseen events being the hardest.
+We observe the best model in each category.
+They obtain weaker performance on unseen splits, as expected, unseen events being the hardest.
 Contextualized models achieve higher margins.
@@
 6.4 Target Entity Visualization
-In table 4, we study an event, Kavanaugh Supreme Court Nomination, using entity target sentiment annotations.
-We identify entities that are discussed by both parties.
-We further separate them into divisive and agreed-upon entities based on expressed sentiments.
-We also show partisan discussed entities for the event.
-This analysis paints a very accurate picture of the discussed event.
+The main goal of this analysis is to demonstrate the usefulness and inspire modeling research in the direction of entity-sentiment-centric view of political events.
+table 4 visualizes one component of how partisan discourse is structured in these events.
+We study Kavanaugh Supreme Court Nomination.
+We identify discussed entities and separate them into divisive and agreed-upon entities.
+This analysis paints an accurate picture of the discussed event.
 We observe that the main entities of Trump, Dr.
@@
 Entities such as Susan Collins and Anita Hill who were vocal mouthpieces of the respective party stances but didn’t directly participate in the event emerge as partisan entities.
-Supreme Court, FBI, and other entities occur in the discourse but only as neutral entities.
-This analysis shows the usefulness of the Target Entity and Sentiment identification in accurately summarizing events.
-Successful models on these tasks could greatly help in understanding political events better.
+Supreme Court, FBI, and other entities occur but only as neutral entities.
 6.5 DCF Context Understanding
-We look into examples that are incorrectly predicted using Wikipedia pages but correctly predicted when using the DCF model.
-We report some of the examples in table 6 in the appendix.
-In examples $11$ & $22$ of Target Entity-Sentiment task, we see that when the entity is not explicitly mentioned in the tweet, the Wiki-Context model fails to identify them as the target entities.
-We posit that while the Wikipedia page of each relevant event will contain these names, explicit modeling of entities in DCF model allows these examples to be correctly classified.
+We look into examples that are incorrectly predicted using Wikipedia pages but correctly predicted by the DCF model in the appendix (table 6).
+In examples $11$ & $22$ of Target Entity-Sentiment task, when the entity is not explicitly mentioned in the tweet, the Wiki-Context model fails to identify them as the targets.
+We posit that while the Wikipedia page of each relevant event will contain these names, explicit modeling of entities in the DCF model allows correct classification.
 Examples 13131-31 - 3 of Vague Text Disambiguation task show that when no clear terms indicate the sentiment towards a view, the Wiki-Context model fails to disambiguate the tweet text.
-Explicit modeling of politician nodes in the DCF model might allow these texts to be correctly disambiguated.
+Explicit modeling of politician nodes seems to help the DCF model.
 7 Conclusion and Future Work
-In this paper, we motivate, define, and operationalize “Social Context Grounding” for political discourse.
-We build two datasets that are useful to evaluate social context grounding in NLP models.
-We experiment with many types of context modeling and benchmark SOTA models.
-We show that explicit modeling of social context outperforms naive concatenation of context and LLMs such as GPT-3.
-Future work includes building datasets for other components of Social Context Grounding and models that account for context in various forms.
+In this paper, we motivate, define, and operationalize ‘Social Context Grounding’ for political text.
+We build two novel datasets to evaluate social context grounding in NLP models that ‘are easy for humans’ when the relevant social context is provided.
+We experiment with many types of contextual models.
+We show that explicit modeling of social context outperforms other models while lacking behind humans.
+Acknowledgements
+We thank Shamik Roy, Nikhil Mehta, and the anonymous reviewers for their vital feedback.
+The project was funded by NSF CAREER award IIS-2048001 and the DARPA CCU Program.
+The contents are those of the author(s) and do not necessarily represent the official views of, nor an endorsement by, DARPA, or the US Government.
 Limitations
@@
 Although we employed human validation at each stage, biases could creep into the datasets.
-We also don’t account for the completeness of out datasets as it is a pioneering work on a new problem.
-Social context is vast and could have myriad of components.
+We also don’t account for the completeness of our datasets as it is a pioneering work on a new problem.
+Social context is vast and could have a myriad of components.
 We only take a step in the direction of social context grounding in this work.
 The performance on these datasets might not indicate full social context understanding but they should help in sparking research in the direction of models that explicitly model such context.
-Although we tuned our prompts a lot, it is possible that better prompts and evolving models might produce better results on the LLM baselines.
+Although we tuned our prompts a lot, better prompts and evolving models might produce better results on the LLM baselines.
 Our qualitative analysis is predicated on a handful of examples.
 They are attempts to interpret the results of large neural models and hence don’t carry as much confidence as our empirical observations.
+We believe the insights from our findings will encourage more research in this area.
+For example, the development of discourse contextualized models that aim to model human-style understanding of background knowledge, emotional intelligence, and societal context understanding is a natural next step of our research.
 Ethics Statement
@@
 It would also be able to ground the text in news articles that expand upon the short texts to provide full context.
-Appendix A Reproducibility
+Appendix A GPT Prompts
+Prompts for Target-Entity Task:
+Event: <event>
+Event background: <background-description>
+Tweet: <tweet-text>
+Author: <author-name>
+Author Party: <party-affiliation>
+Author background: <first two sentences of author-wiki-page>
+Target Entity: <entity-name>
+Entity background: <first two sentences of entity-wiki-page>
+Task: Identify if the given entity is a target of the tweet. A target entity is defined as an entity that would be present in the full unambiguous explanation of the tweet.
+Is the given entity a target entity of the tweet? Answer yes or no.
+Prompts for Target-Sentiment Task:
+Event: <event>
+Event background: <background-description>
+Tweet: <tweet-text>
+Author: <author-name>
+Author Party: <party-affiliation>
+Author background: <first two sentences of author-wiki-page>
+Target Entity: <entity-name>
+Entity background: <first two sentences of entity-wiki-page>
+Task: Identify the sentiment of the tweet towards the given target entity. Consider that the tweet is ambiguous and the entity might be implied without being explicitly mentioned.
+What is the sentiment of the tweet towards the target entity? Answer with positive, negative, or neutral.
+Prompts for Vague Text Task:
+Event: <event>
+Event background: <background-description>
+Vague message: <vague-text>
+Author Party: <party-affiliation>
+Author background: <first two sentences of party-wiki-page>
+Task: Given the event, vague message, and party affiliation of the author, explain unambiguously the intended meaning of the vague message.
+Generate an unambiguous explanation for the vague message given the party affiliation of the author and the event in context.
+Appendix B Reproducibility
 We use the HuggingFace Transformers (Wolf et al., 2020) library for PLMs.
@@
 We experiment with a random seed value set to {13131313}.
-We report the results of the $33$ fold cross-validation.
-We report only the mean on all the cross-fold validation results for clarity.
-All our code, datasets, and result logs will be released publicly upon acceptance.
-We experiment with $33$, $55$, and $1010$ fold cross-validation.
-As the results on the development data are almost identical, we report the results of $33$ fold cross-validation in all our experiments.
-Appendix B Error Analysis
+We set CUBLAS environment variables for reproducibility.
+All our code, datasets, and result logs are released publicly.
+Appendix C Error Analysis
 Target Entity and Sentiment Task Vague Text Disambiguation Task
@@
 Table 6: Examples where baseline model fails but DCF works
-Appendix C Annotation Interfaces
+Appendix D Annotation Interfaces
 Figure 2: An example of Tweet Target Entity and Sentiment Annotation GUI